Can I please get some help with this question?
Time on a treadmill. After 2 minutes on a treadmill Jenny has a heart rate of 82. After 3 mins she has a heart rate of 86. Assume that there is a linear equation that gives her heart rate h in terms of time on the tread mill t. Find the equation and use it to predict her heart rate after 10 mins on the treadmill.

Answer by venugopalramana(3286)

Time on a treadmill. After 2 minutes on a treadmill Jenny has a heart rate of 82. After 3 mins she has a heart rate of 86. Assume that there is a linear equation that gives her heart rate h in terms of time on the tread mill t. Find the equation and use it to predict her heart rate after 10 mins on the treadmill.
AT T=2...H=82
AT T=3....H=86
INCREAE IN TIME =3-2=1
INCREASE IN HEART RATE IN THIS PERIOD = 86-82=4
RATE OF INCREASE PER UNIT TIME = 4/1=4
IN LINEAR RELATION THIS REMAINS CONSTANT.
SO TAKING BASE VALUE AS T=2 AND H=82..WE GET FOR ANT TIME T AND RATE H
INCREASE IN TIME =T-2
INCREASE IN RATE=H-82
RATE OF INCREASE =(H-82)/(T-2)=4
H-82=4(T-2)=4T-8
H-4T=82-8=74
SO THE EQN. IS
H-4T=74
HENCE AT T=10 WE GET
H=74+4*10=114